OpenStudy (anonymous):

How do you simplify (-3-2i)^2

OpenStudy (anonymous):

5+12 i

OpenStudy (stapeleliz):

Use the exact same process as you would for multiplying out (3 - 2x)^2. The only difference is that you'll be able to simplify i^2 to -1, where x^2 would just stay x^2.

OpenStudy (anonymous):

is this a complex number? the expansion looks like this (-3-2i)(-3-2i) first term: -2i*-2i=4i^2=4(-1) second term: -2i*-3+-2i*-3=12i third term: -3*-3=9 final expansion=-4+9+12i=5+12i